Tohban Report 2014-04-16

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 9 Apr 2014
End Date: 16 Apr 2014
Tohban: Mitsuo Oka
Tohban email:
Next Tohban: Lindsay Glesener
List all reports


Solar Activity

The solar activity was relatively low throughout the week although there were numerous C-class flares. There are currently 9 active regions on the disk and STEREO B shows no significantly bright region that would rotate onto the disk within a few days.

How many GOES flares occurred?

 Flares above B, C, M, X class were     13    34     0     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      6    19     0     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      2     8     0     0

There were RHESSI flares/GOES flares 133 / 47 over the time range 08-Apr-14 15-Apr-14

Memory Management

The SSR fill level had been relatively high (20-40%) throughout the week.

Spacecraft Status

The cold plates, 1 and 2 were at 118.2 and 116.7 K, respectively, on 2014-04-16. The spin rate is currently 14.6 and we need spin-up soon.

Data Gaps

No data gap

Detector issues

Many threshold/HV changes were made in G4 rear, G6 front and G8 rear, because their livetime kept decreasing. The G9 change was made to try to restore recorded events.

Other notes

The thin attenuator is stuck in again, starting at the end of April 13 (Noticed by Brian). The attenuator logic currently uses G1, G4, G5, and a decision has been made at the Ops meeting that the logic will be changed to use G1, G2, G5, and G7. Tohbans should continue to be watchful for stuck attenuators because G5's resting deadtime looks to be creeping up. Once we can no longer use G5, we will need to start doubling up on detectors.

Spacecraft Management

Decimation Active/Vigorous
HLAT Decimation Rear decimation weight 6, no front decimation
Night time data (fronts) +/- 4 minutes
Night time data (rears) +/- 4 minutes
Require extra passes? Yes, for ADP data
Requirement for moving pointer? No
Attenuator operation Normal
Detector problems? See motes above.
Personal tools